OpenStudy (anonymous):

What is the equation of a line passing through (-6, 5) and having a slope of 1/3 ?

OpenStudy (anonymous):

you could use y=mx+b

OpenStudy (anonymous):

m is the slope plug it in

OpenStudy (anonymous):

y=1/3x+b but now how do I find b?

OpenStudy (anonymous):

To find b, think about what your (x,y) point means: (-6,5). When x of the line is -6, y of the line must be 5. Because you said the line passes through this point, right? Now, look at our line's equation so far: . b is what we want, the 1/3 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the the point (-6,5). So, why not plug in for x the number -6 and for y the number 5? This will allow us to solve for b for the particular line that passes through the point you gave!. (-6,5). y=mx+b or 5=1/3 × -6+b, or solving for b: b=5-(1/3)(-6). b=7.

Nnesha (nnesha):

use that order pair where (x,y) values are replace x and y by their valeus and solve for b

Nnesha (nnesha):

values*

Nnesha (nnesha):